OpenStudy (lgbasallote):

how to get R in \[\frac{30}{e} = e^{-5/R} (-\frac{250}{R} + \frac{100}{R^2})\]

OpenStudy (blockcolder):

Graph them, I suppose...

OpenStudy (cherylim23):

This is a transcendental function with a linear function... It wouldn't be possible to solve unless you use Newton's Approximation. http://www.math.brown.edu/UTRA/linapprox.html

OpenStudy (lgbasallote):

lol isnt this just algebra

OpenStudy (lgbasallote):

i just need to get R for my diff eq thingy

OpenStudy (lgbasallote):

ugh what did i do wrong :/

OpenStudy (anonymous):

dyslexic i am today... sorry...

OpenStudy (lgbasallote):

what is the solution for \[\frac{dq}{dt} + \frac{q}{0.02R} = \frac{100}{R}\]

OpenStudy (valpey):

\[q= Ae^{\frac{-50}{R}t}+2\]

OpenStudy (valpey):

Unless q is just a function of R.

OpenStudy (lgbasallote):

how did you get this?

OpenStudy (valpey):

Pretty much the same thing we did before. I reasoned, If q were constant, what would it have to be to get 100/R :: answer 2.

OpenStudy (lgbasallote):

what i got was \[\LARGE q = \frac 2R + Ae^{-\frac{t}{0.02R}}\]

OpenStudy (valpey):

Of course 50t/R = t/0.02R so we got that part the same. But I think you have q/.02R in your equation so if you plug in your answer you will have a R^2 in the denominator.

OpenStudy (valpey):

\[\frac{2}{.02R}=\frac{100}{R}\] \[\frac{2/R}{.02R}<>\frac{100}{R}\]

OpenStudy (valpey):

Essentially, set A=0 and see if it solves the ODE.

OpenStudy (lgbasallote):

oh lol R cancels out *facepalm*

OpenStudy (lgbasallote):

wait lemme redo this

OpenStudy (lgbasallote):

ugh my end answer is still wrong >.<

OpenStudy (lgbasallote):

okay...how to solve for R in \[\large \frac{30}{e} = -\frac{3e^{-\frac 5R}}{R}\]

OpenStudy (lgbasallote):

this is where it came from...maybe i just did something wrong http://openstudy.com/updates/4ffd2fc8e4b00c7a70c5dc86